# This is a preprint submitted to: Computer Methods in Applied Mechanics and Engineering

### v1

@@ 2.2 Physics-informed parametric model embedding
$\mathbf{P}_{I}=\left[\begin{array}[]{c}\mathbf{D}\\ \mathbf{U}\\ \mathbf{F}\\ \mathbf{C}\end{array}\right]\qquad\mathrm{with}\qquad\mathbf{F}=\left[\begin{array}[]{ccc}|&\vdots&|\\ \hat{{\mathbf{f}}}_{j}^{(1)}&\dots&\hat{{\mathbf{f}}}_{j}^{(S)}\\ |&\vdots&|\\ \end{array}\right]\qquad\mathrm{and}\qquad\mathbf{C}=\left[\begin{array}[]{ccc}|&\vdots&|\\ \hat{c}_{j}^{(1)}&\dots&\hat{c}_{j}^{(S)}\\ |&\vdots&|\\ \end{array}\right]$ (16)
@@
${\mathbf{W}}_{I}=\left[\begin{array}[]{cccc}\mathbf{W}_{\mathbf{d}}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{W}_{\mathbf{u}}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{W}_{\mathbf{f}}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{W}_{c}\\ \end{array}\right]=\mathrm{diag}\Bigl{(}\mathbf{W}_{\mathbf{d}},\underbrace{\mathbf{0}}_{\mathbf{u}},\mathbf{W}_{\mathbf{f}},\mathbf{W}_{c}\Bigr{)},$ (17)
where $\mathbf{W}_{\mathbf{d}}$ is the block-diagonal matrix weighting the geometric entries (as in standard PME), $\mathbf{W}_{\mathbf{u}}$ the one for design variables, $\mathbf{W}_{\mathbf{f}}$ for distributed physical data $\mathbf{F}$, and $\mathbf{W}_{c}$ for lumped scalars $\mathbf{C}$. Each block has to be set up to normalize its respective data by the inverse of the estimated variance. In practice, for a block associated with a data vector $y\in\{{\mathbf{d}},{\mathbf{f}},c\}$, one calculates
$\mathrm{Var}(y)=\sum_{j=1}^{S}(y_{j}-\langle y\rangle)^{2}/S,$ (18)
@@
If the block contains multiple components (e.g., multi-dimensional fields or multiple operating conditions), each row or column of that block can be assigned its own weight $\rho_{i}$. Doing so ensures that geometric modifications and physical data all contribute comparably to the subsequent PCA procedure—i.e., no portion of the data dominates simply because it exhibits a larger raw variance.
@@
$\mathbf{A}_{I}=\frac{1}{S}\mathbf{P}_{I}\mathbf{P}_{I}^{\mathsf{T}},\qquad\mathbf{G}_{I}=\left[\begin{array}[]{cccc}\mathbf{G}_{\mathbf{d}}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{I}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{G}_{\mathbf{f}}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{I}\\ \end{array}\right],\qquad\mathbf{Z}_{I}=\left[{\mathbf{z}}_{I,1}\dots\mathbf{z}_{I,S}\right],\qquad{\mathbf{z}}_{I,k}=\left[\begin{array}[]{c}\mathbf{q}_{k}\\ {\mathbf{v}}_{k}\\ \boldsymbol{\phi}_{k}\\ \boldsymbol{\pi}_{k}\end{array}\right]_{I}$ (21)
@@ 2.3 Physics-driven parametric model embedding
which assigns, as per standard PME and PI-PME, zero weight to the rows corresponding to $\mathbf{U}$, while normalizing the distributed physical fields $\mathbf{F}$ and the lumped scalars $\mathbf{C}$ by the inverse of their respective variances, as per PI-PME (see, eqs. 18 and 19).
@@
${\mathbf{A}}_{D}\;{\mathbf{G}}_{D}\;{\mathbf{W}}_{D}\;{\mathbf{Z}}_{D}\;=\;{\mathbf{Z}}_{D}\;\boldsymbol{\Lambda}_{D},\quad\text{where}\quad{\mathbf{A}}_{D}\;=\;\frac{1}{S}\,\mathbf{P}_{D}\,\mathbf{P}_{D}^{\mathsf{T}},$
@@ 3.1 Test case 1: RAE-2822 airfoil
The first test case concerns the design optimization of the classical RAE-2822 airfoil (see Fig. 2, bottom left). The design space, defined within the activities of the NATO-AVT-331 [33], includes $M=20$ design variables; each variable is associated with a different shape function (3 polynomials, 6 Hicks–Henne bumps, and 1 Wagner function [34], see Fig. 2, right), acting either on the upper or lower surface of the airfoil [35]. An in-house code (WG2AER, developed at CIRA) parameterizes the airfoil as a linear combination of the parent geometry ${\mathbf{g}}({\boldsymbol{\xi}})$ and the modification functions ${\boldsymbol{\delta}}$.
@@ 3.2 Test case 2: autonomous underwater glider
The second test case addresses the design optimization of a bio-inspired autonomous underwater glider (AUG) with a manta-like shape, that has been selected as a test case for the activity of the NATO-AVT-404 on “Enhanced Design Processes of Military Vehicles through Machine Learning Methods”. The geometry is constructed as a continuous wing using a section-wise scheme and is by design spanwise symmetric; only half of the body is parameterized and then mirrored at the root section. As illustrated in Fig. 3, the model comprises three zones: the center body, the transition region, and the outer wing. Four sections define the half-span: (i) the root (section 1), (ii) the end of the main body (section 2), (iii) the end of the transition (section 3), and (iv) the tip (section 4). Each section is fully determined by ten variables: four parameters for the section’s NACA 4-digit airfoil (maximum camber $m$, maximum camber position $p$, thickness ratio $t$, chord $c$); three parameters for positioning the section’s leading edge $(x_{0},z_{0},s_{0})$; three parameters defining the rotation angles (pitch $\vartheta$, roll $\phi$, yaw $\psi$), applied sequentially about the leading edge. Section 1 retains only 2 degrees of freedom ($t$ and $c$), while all other parameters are fixed. Consequently, the complete parametric model includes $M=32$ design variables. The manta-like geometry is constructed via the OpenCASCADE CAD kernel [37] and meshed with Gmsh v4.12.1 [38].
@@ 3.3 Test case 3: ship propeller
(a) Radial distributions (chord, pitch, rake, and maximum sectional camber, from left to right)
(b) Sectional distributions (chordwise thickness and camber line, left and right)
Figure 5: Test case 3: distributions of the geometrical features of the propeller
@@
For this particular problem, the blade geometry is parametrized through B-Spline control polygons defining radial distributions of non-dimensional chord ($c/D$, parameter name $c$), pitch ($P/D$, parameter name $pd$), rake ($rake/D$, parameter name $r$), and maximum sectional camber ($f_{max}/c$, parameter name $f$) of Fig. 5a. The shape of sectional hydrofoils is included as well in the parametric model by describing, again by using appropriate control polygons (see Fig. 5b), the non-dimensional chordwise sectional thickness ($t/2t_{max}$, parameter name $th$) and camber ($f/f_{max}$, parameter name $fh$) distributions of the blade profile which, radial section per radial section, are finally scaled with the corresponding maximum values. Given some constraints (the radial position of control points at root and tip is fixed, leading- and trailing-edge points are always given, and the chord at the tip never changes to comply with some limits of the flow solver adopted for the hydrodynamic characterization of performances), control points, sequentially numbered from root to tip (or from leading- to trailing-edge), are free to move within assigned ranges. Subscript “x” indicates the radial, or chordwise, modification of the control point position, and subscript “y” refers instead to changes in the quantity the control point describes. This parametrization leads to a total of $M=38$ design variables (5 for the chord, 8 for the maximum sectional camber and pitch each, 6 for the rake, 6 for the sectional camber line, and 5 for the sectional thickness non-dimensional distribution) which describe entirely the blade shape with the exception of the skew that is maintained unchanged and identical to that of the reference propeller. For the hydrodynamic analyses, a low-fidelity boundary element method (BEM) is employed. The BEM code [42] was developed at the University of Genoa since early 2000 for analysis and design-by-optimization purposes. It is a Morino, Dirichlet-type boundary condition implementation of panel methods for incompressible, potential flow solution, which makes use of the “key blade” approach to deal with stationary or unsteady problems [43]. It includes a cavitation model (the sheet cavity model at leading edge and midchord, both on suction and pressure side proposed in [44]), wake alignment capabilities and the iterative Kutta condition. For current analyses, the “key blade” is discretized with 1250 hyperboloidal panels (Ł= 1326 nodes). The collection of physical information was carried out by considering several equivalent steady-state operating conditions representative of the most critical conditions encountered by the blade during a revolution, given the design of the reference propeller, which was developed to operate in the spatially non-uniform wake of the ship shown in Fig. 6.
Figure 6: Non-uniform inflow (ship wake) to the propeller, seen from aft.
@@ 3.4 Test case 4: 5415 destroyer-type vessel
(a) TC#1: RAE-2822
(b) TC#2: AUG
(c) TC#3: Propeller
(d) TC#4: 5415
Figure 8: Variance resolved as a function of the reduced design variables
@@ 4 Results
To train the standard PME and its physics-informed extensions (PI-PME and PD-PME), an initial set of $S=16385$ Sobol samples was generated for each test case. In addition to covering the geometric design space, physical simulations were performed at these samples whenever possible, providing distributed or lumped physical data. Any geometries that produced unfeasible outcomes (e.g., degenerate shapes, solver divergence, or NaN outputs) were discarded. This filtering process alone already serves as a preliminary assessment of the chosen parameterization, as it highlights whether certain geometric variations are physically inadmissible or numerically unstable.
An additional layer of screening was then applied to exclude extreme outliers. Specifically, all geometric–physical samples were retained only if the corresponding physical outputs fell within the ranges $[Q1-3IQR,Q3+3IQR]$, where $Q1$ and $Q3$ are the first and third quartiles, respectively, and $IQR=Q3-Q1$. This so-called extreme outlier rule was chosen to be sufficiently permissive, while remaining distribution-agnostic. As a result, only physically credible shapes and simulations populate the final data set used for dimensionality reduction.
@@ 3 Propeller 13,615 38 8 (79%) 12 (68%) 8 (79%)
PME, PI-PME, and PD-PME curves in the RAE airfoil problem lie close together (see Fig. 8a); each method needs a comparable number of modes ($N=6$ or 7) to reach 95% or 99% variance. This suggests that geometric changes coincide strongly with variations in the pressure distribution and integrated aerodynamic coefficients. Figure 9 (left) shows a pretty similar embedding with eigenvectors almost coincident for PME, PI-PME, and PD-PME (see first and third mode), in particular, it can be seen how the first mode is mainly participated by $u_{2}$, $u_{4}$, and $u_{6}$, that correspond to the polynomial shape modification function applied to the lower side of the airfoil, highlighting also how the lower side modification is the major region affecting the aerodynamic performance. This is further confirmed by Figure 9 (right) where is shown that geometry (${\mathbf{d}}$) and physical parameters ($\mathbf{c}_{p}$, $c_{L}$, $c_{D}$, and $c_{M}$) share most of the first modes: apart from mode 2, that is mainly participated by the drag coefficient, no distinct pure physics or pure geometry mode dominates the decomposition. Because airfoil shape changes strongly affect the aerodynamic response under these conditions, the geometry and physics are highly correlated. Hence, adding physical data (PI-PME) does not inflate the necessary dimensionality, and removing geometry altogether (PD-PME) does not drastically reduce the dimension beyond PME, meaning that both geometry and physics point in similar directions in the design space.
@@
For the glider test case, the PME rapidly achieves a high fraction of geometric variance ($N=8$ modes for 95%), whereas the PI-PME requires more modes ($N=11$). In contrast, PD-PME dips below the PME line, reaching 95% in fewer modes ($N=5$), as shown in Fig. 8b. Figure 10, shows, on the left, that while for PME the highest variance (retained by the first mode) is mainly due to the span ($s_{0,4}$) of the tip section, the variance of PI-PME and PD-PME are mainly affected by the maximum camber and the pitch angle of the end of the transition from the center body to the outer wing ($m_{3}$ and $\vartheta_{3}$). As shown in Fig. 10 (right) the first modes appear “decoupled” (i.e., geometry heavy or physics heavy): for both PI-PME and PD-PME the first mode is mainly participated by global forces ($L$ and $D$), the second mainly by the pressure distribution $\mathbf{c}_{p}$, and the third mainly by the geometry (for PI-PME only). There is a partial uncorrelation between geometry and physics: large shape modifications do not necessarily produce proportional changes in lift or drag. Hence, PI-PME must retain more overall modes to capture both geometry-dominated and physics-dominated directions. Meanwhile, if the goal is purely performance-driven, PD-PME focuses on those physical variations alone, achieving a stronger reduction with fewer modes.
@@
(a) TC#1: RAE-2822
(b) TC#2: AUG
(c) TC#3: Propeller
(d) TC#4: 5415
Figure 13: Contour maps of lumped physical information as a function of the first two reduced design variables ${\mathbf{x}}$
(a) RAE-2822
(b) AUG
(c) Propeller
(d) DTMB-5415
Figure 14: Scatter plot of lumped physical parameters usable as objective functions in a multi-objective problem

### v2

@@ 2.2 Physics-informed parametric model embedding
$\mathbf{P}_{I}=\left[\begin{array}[]{c}\mathbf{D}\\ \mathbf{U}\\ \mathbf{F}\\ \mathbf{C}\end{array}\right]$ (19)
with
$\mathbf{F}=\left[\begin{array}[]{ccc}|&\vdots&|\\ \hat{{\mathbf{f}}}_{j}^{(1)}&\dots&\hat{{\mathbf{f}}}_{j}^{(S)}\\ |&\vdots&|\\ \end{array}\right]$ (20)
and
$\mathbf{C}=\left[\begin{array}[]{ccc}|&\vdots&|\\ \hat{c}_{j}^{(1)}&\dots&\hat{c}_{j}^{(S)}\\ |&\vdots&|\\ \end{array}\right]$ (21)
@@
${\mathbf{W}}_{I}$ $=\left[\begin{array}[]{cccc}\mathbf{W}_{\mathbf{d}}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{W}_{\mathbf{u}}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{W}_{\mathbf{f}}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{W}_{c}\\ \end{array}\right]$ (22)
$=\mathrm{diag}\Bigl{(}\mathbf{W}_{\mathbf{d}},\underbrace{\mathbf{0}}_{\mathbf{u}},\mathbf{W}_{\mathbf{f}},\mathbf{W}_{c}\Bigr{)},$
where $\mathbf{W}_{\mathbf{d}}$ is the block-diagonal matrix weighting the geometric entries (as in standard PME), $\mathbf{W}_{\mathbf{u}}$ the one for design variables, $\mathbf{W}_{\mathbf{f}}$ for distributed physical data $\mathbf{F}$, and $\mathbf{W}_{c}$ for lumped scalars $\mathbf{C}$.
Each block has to be set up to normalize its respective data by the inverse of the estimated variance. In practice, for a block associated with a data vector $y\in\{{\mathbf{d}},{\mathbf{f}},c\}$, one calculates
$\mathrm{Var}(y)=\frac{1}{S}\sum_{j=1}^{S}(y_{j}-\langle y\rangle)^{2},$ (23)
@@
If the block contains multiple components (e.g., multi-dimensional fields or multiple operating conditions), each row or column of that block can be assigned its own weight $\rho_{i}$. This normalization strategy ensures that both geometric modifications and physical observables contribute on a comparable scale to the PCA procedure—preventing any data block from dominating the embedding solely due to differences in units, magnitude, or raw variance.
@@
$\mathbf{A}_{I}=\frac{1}{S}\mathbf{P}_{I}\mathbf{P}_{I}^{\mathsf{T}},\qquad\mathbf{G}_{I}=\left[\begin{array}[]{cccc}\mathbf{G}_{\mathbf{d}}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{I}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{G}_{\mathbf{f}}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{I}\\ \end{array}\right],$ (26)
and
$\mathbf{Z}_{I}=\left[{\mathbf{z}}_{I,1}\dots\mathbf{z}_{I,S}\right],\qquad{\mathbf{z}}_{I,k}=\left[\begin{array}[]{c}\mathbf{q}_{k}\\ {\mathbf{v}}_{k}\\ \boldsymbol{\phi}_{k}\\ \boldsymbol{\pi}_{k}\end{array}\right]_{I}$ (27)
@@ 2.3 Physics-driven parametric model embedding
which assigns, as per standard PME and PI-PME, zero weight to the rows corresponding to $\mathbf{U}$, while normalizing the distributed physical fields $\mathbf{F}$ and the lumped scalars $\mathbf{C}$ by the inverse of their respective variances, as per PI-PME (see, eqs. 23 and 24).
@@
${\mathbf{A}}_{D}\;{\mathbf{G}}_{D}\;{\mathbf{W}}_{D}\;{\mathbf{Z}}_{D}\;=\;{\mathbf{Z}}_{D}\;\boldsymbol{\Lambda}_{D},$ (28)
where
${\mathbf{A}}_{D}\;=\;\frac{1}{S}\,\mathbf{P}_{D}\,\mathbf{P}_{D}^{\mathsf{T}},$ (29)
@@ 3.1 Test case 1: RAE-2822 airfoil
The first test case concerns the design optimization of the classical RAE-2822 airfoil (see Fig. 2, bottom left). The design space, defined within the activities of the NATO-AVT-331 Research Task Group [35], includes $M=20$ design variables; each variable is associated with a different shape function (3 polynomials, 6 Hicks–Henne bumps, and 1 Wagner function [36], see Fig. 2, right), acting either on the upper or lower surface of the airfoil [37]. An in-house code (WG2AER, developed at CIRA) parameterizes the airfoil as a linear combination of the parent geometry ${\mathbf{g}}({\boldsymbol{\xi}})$ and the modification functions ${\boldsymbol{\delta}}$.
@@ 3.2 Test case 2: autonomous underwater glider
The second test case addresses the design optimization of a bio-inspired autonomous underwater glider (AUG) with a manta-like shape, that has been selected as a test case for the NATO-AVT-404 Research Task Group on “Enhanced Design Processes of Military Vehicles through Machine Learning Methods”. The geometry is constructed as a continuous wing using a section-wise scheme and is by design spanwise symmetric; only half of the body is parameterized and then mirrored at the root section. As illustrated in Fig. 3, the model comprises three zones: the center body, the transition region, and the outer wing. Four sections define the half-span: (i) the root (section 1), (ii) the end of the main body (section 2), (iii) the end of the transition (section 3), and (iv) the tip (section 4). Each section is fully determined by ten variables: four parameters for the section’s NACA 4-digit airfoil (maximum camber $m$, maximum camber position $p$, thickness ratio $t$, chord $c$); three parameters for positioning the section’s leading edge $(x_{0},z_{0},s_{0})$; three parameters defining the rotation angles (pitch $\vartheta$, roll $\phi$, yaw $\psi$), applied sequentially about the leading edge. Section 1 retains only 2 degrees of freedom ($t$ and $c$), while all other parameters are fixed. Consequently, the complete parametric model includes $M=32$ design variables. The manta-like geometry is constructed via the OpenCASCADE CAD kernel [39] and meshed with Gmsh v4.12.1 [40].
@@ 3.3 Test case 3: ship propeller
Figure 6: Non-uniform inflow (ship wake) to the propeller, seen from aft
For this particular problem, the blade geometry is parametrized through B-Spline control polygons defining radial distributions of non-dimensional chord ($c/D$, parameter name $c$), pitch ($P/D$, parameter name $pd$), rake ($rake/D$, parameter name $r$), and maximum sectional camber ($f_{max}/c$, parameter name $f$) of Fig. LABEL:fig:prop_sezionia. The shape of sectional hydrofoils is included as well in the parametric model by describing, again by using appropriate control polygons (see Fig. LABEL:fig:prop_sezionib), the non-dimensional chordwise sectional thickness ($t/2t_{max}$, parameter name $th$) and camber ($f/f_{max}$, parameter name $fh$) distributions of the blade profile which, radial section per radial section, are finally scaled with the corresponding maximum values. Given some constraints (the radial position of control points at root and tip is fixed, leading- and trailing-edge points are always given, and the chord at the tip never changes to comply with some limits of the flow solver adopted for the hydrodynamic characterization of performances), control points, sequentially numbered from root to tip (or from leading- to trailing-edge), are free to move within assigned ranges. Subscript “x” indicates the radial, or chordwise, modification of the control point position, and subscript “y” refers instead to changes in the quantity the control point describes. This parametrization leads to a total of $M=38$ design variables (5 for the chord, 8 for the maximum sectional camber and pitch each, 6 for the rake, 6 for the sectional camber line, and 5 for the sectional thickness non-dimensional distribution) which describe entirely the blade shape with the exception of the skew that is maintained unchanged and identical to that of the reference propeller. For the hydrodynamic analyses, a low-fidelity boundary element method (BEM) is employed. The BEM code [44] was developed at the University of Genoa since early 2000 for analysis and design-by-optimization purposes. It is a Morino, Dirichlet-type boundary condition implementation of panel methods for incompressible, potential flow solution, which makes use of the “key blade” approach to deal with stationary or unsteady problems [45]. It includes a cavitation model (the sheet cavity model at leading edge and midchord, both on suction and pressure side proposed in [46]), wake alignment capabilities and the iterative Kutta condition. For current analyses, the “key blade” is discretized with 1250 hyperboloidal panels (Ł= 1326 nodes). The collection of physical information was carried out by considering several equivalent steady-state operating conditions representative of the most critical conditions encountered by the blade during a revolution, given the design of the reference propeller, which was developed to operate in the spatially non-uniform wake of the ship shown in Fig. 6.
@@ 4 Results
To train the standard PME and its physics-informed extensions (PI-PME and PD-PME), an initial set of $S=16385$ samples was generated for each test case. To ensure a uniform and representative coverage of the design space, the initial dataset is generated using a Sobol low-discrepancy sequence. This approach provides better space-filling properties than purely random sampling and avoids clustering, which is critical to reliably constructing the reduced manifold. In addition to covering the geometric design space, physical simulations were performed at these samples whenever possible, providing distributed or lumped physical data. Any geometries that produced unfeasible outcomes (e.g., degenerate shapes, solver divergence, or NaN outputs) were discarded. This filtering process alone already serves as a preliminary assessment of the chosen parameterization, as it highlights whether certain geometric variations are physically inadmissible or numerically unstable.
An additional layer of screening was then applied to exclude extreme outliers. Specifically, all geometric–physical samples were retained only if the corresponding physical outputs fell within the ranges $[Q1-3IQR,Q3+3IQR]$, where $Q1$ and $Q3$ are the first and third quartiles, respectively, and $IQR=Q3-Q1$. This distribution-agnostic rule was selected to balance permissiveness and robustness: it allows broad variability across designs while excluding pathological outliers. As a result, only physically credible shapes and simulations populate the final data set used for dimensionality reduction. The filtering strategy is intentionally cautious, aiming to eliminate anomalous or ill-posed configurations that could contaminate the reduced basis with non-representative trends. It is finally worth noting that, while the reduced space may still contain directions that could lead to poorly behaved geometries, standard constraint-handling techniques in the subsequent optimization process can be employed to reject such configurations if encountered.
@@ 3 Propeller 13,615 38 8 (79%) 12 (68%) 8 (79%)
PME, PI-PME, and PD-PME curves in the RAE airfoil problem lie close together (see Fig. LABEL:fig:eigsumsa); each method needs a comparable number of modes ($N=6$ or 7) to reach 95% or 99% variance. This suggests that geometric changes coincide strongly with variations in the pressure distribution and integrated aerodynamic coefficients. Figure 11 (left) shows a pretty similar embedding with eigenvectors almost coincident for PME, PI-PME, and PD-PME (see first and third mode), in particular, it can be seen how the first mode is mainly participated by $u_{2}$, $u_{4}$, and $u_{6}$, that correspond to the polynomial shape modification function applied to the lower side of the airfoil, highlighting also how the lower side modification is the major region affecting the aerodynamic performance. This is further confirmed by Figure 11 (right) where is shown that geometry (${\mathbf{d}}$) and physical parameters ($\mathbf{c}_{p}$, $c_{L}$, $c_{D}$, and $c_{M}$) share most of the first modes: apart from mode 2, that is mainly participated by the drag coefficient, no distinct pure physics or pure geometry mode dominates the decomposition. Because airfoil shape changes strongly affect the aerodynamic response under these conditions, the geometry and physics are highly correlated. Hence, adding physical data (PI-PME) does not inflate the necessary dimensionality, and removing geometry altogether (PD-PME) does not drastically reduce the dimension beyond PME, meaning that both geometry and physics point in similar directions in the design space.
@@
For the glider test case, the PME rapidly achieves a high fraction of geometric variance ($N=8$ modes for 95%), whereas the PI-PME requires more modes ($N=11$). In contrast, PD-PME dips below the PME line, reaching 95% in fewer modes ($N=5$), as shown in Fig. LABEL:fig:eigsumsb. Figure 12, shows, on the left, that while for PME the highest variance (retained by the first mode) is mainly due to the span ($s_{0,4}$) of the tip section, the variance of PI-PME and PD-PME are mainly affected by the maximum camber and the pitch angle of the end of the transition from the center body to the outer wing ($m_{3}$ and $\vartheta_{3}$). As shown in Fig. 12 (right) the first modes appear “decoupled” (i.e., geometry heavy or physics heavy): for both PI-PME and PD-PME the first mode is mainly participated by global forces ($L$ and $D$), the second mainly by the pressure distribution $\mathbf{c}_{p}$, and the third mainly by the geometry (for PI-PME only). There is a partial uncorrelation between geometry and physics: large shape modifications do not necessarily produce proportional changes in lift or drag. Hence, PI-PME must retain more overall modes to capture both geometry-dominated and physics-dominated directions. Meanwhile, if the goal is purely performance-driven, PD-PME focuses on those physical variations alone, achieving a stronger reduction with fewer modes.